OpenStudy (brightemmerald):

Evaluate the function below for x = 2. F(x) = x^4 + 6x^3 + 3

OpenStudy (lonelyy__):

Hold on I'll help but I need to solve it out

OpenStudy (lonelyy__):

Do you have answer choices?

OpenStudy (shadowlegendx):

\[F(2) = 2^4 + 6(2)^3 + 3\] \[F(2) = 16 + 6(8) + 3\] \[F(2) = 16 + 48 + 3\] \[F(2) = 67\]

OpenStudy (lonelyy__):

Yes he is right.

OpenStudy (brightemmerald):

Thank you!